OpenStudy (rayman5766):

What is the solution to the equation? b + 0.8 = –1

OpenStudy (rayman5766):

i put -9 as my answer

OpenStudy (anonymous):

wrong

OpenStudy (one098):

subtract -1 and 0.8.

OpenStudy (anonymous):

its -1.8

OpenStudy (gorv):

subtract 0.8 from both side

OpenStudy (rayman5766):

yes i did that i got -1.8 are -9

OpenStudy (rayman5766):

Use a transformation to solve the equation. z – (–0.3) = –2.4 A. z = –2.7 B. z = –2.1 C. z = 2.1 D. z = 2.7

OpenStudy (anonymous):

NO ITS b=-1.8+0.8=-1

OpenStudy (rayman5766):

i put in d 2.7